An angry construction worker throws his wrench downward from a height of 128 feet with an initial velocity of 32 feet per second. The height of the wrench above the ground after seconds is given by S(t) = -16y^2 -32t +128.
a) What is the height of the wrench after 1 second?
b) How long does it take for the wrench to reach the ground?

Answer by lefty4ever26(59)

S(t)=height. t = time in seconds. We want to know height at 1 second. SO plug in 1 for t. and solve for S(t).
S(1)=-16(1)^2-32(1)+128=80. SO at 1 second the wrench is at a height of 80ft.
The wrench will be at the ground at 0 ft. So set the equation equal to 0 and solve for t.
S(t)=-16t^2-32t+128=0
-16(t^2+2t-8)=0
-16(t-2)(t+4)=0
t=2, t=-4 Time can't be negative...so at t=2 seconds the wrench has reached the ground.